OpenStudy (anonymous):

when 2 vectors are parallel what we have ? I mean vector AB//BC= ?

OpenStudy (amistre64):

we have 2 vectors the differ by at most a scalar

OpenStudy (anonymous):

AB*BC=0 ?

OpenStudy (amistre64):

does dotting parallel vectors produce a 0? they aint orthogonal that I can recall

OpenStudy (amistre64):

0,1 0,1 ---- 0+1 = 1 0,1 0,-1 ---- 0-1 = -1

OpenStudy (amistre64):

3 , 4 -3,-4 ----- -9-12 = -18

OpenStudy (anonymous):

AB+BC=-1 ?

OpenStudy (amistre64):

adding vectors produces another vector, not a scalar

OpenStudy (amistre64):

youre gonna have to provide more verbiage if I am to understand what it is your asking :)

OpenStudy (amistre64):

that and 4.-4 = -16 ... not -12 lol

OpenStudy (anonymous):

I have the vectors AB(2,6) and BC(2K-2 , 5K-4) And I need to prove that those vectors are parallel. How I will do that?

OpenStudy (anonymous):

I mean,I need to find fot wich k those vectors are parallel.

OpenStudy (amistre64):

when they differ by only a scalar; by a multiplicative factor. 2 = 2k-2 and 6 = 5k-4

OpenStudy (anonymous):

You could test values of k--

OpenStudy (amistre64):

4 = 2k; k=2 10 = 5k; k=2 right? so when k=2; that is one possible option for parallel

OpenStudy (anonymous):

then we have AB//BC <=> AB = BC

OpenStudy (amistre64):

when ab = bc; they are parallel; but thats not the only case ...

OpenStudy (amistre64):

if you can determine the angle that the vectors span from each other; then you can prove if they are parallel or not as well :) 180 and 0 are parallel vectors

OpenStudy (anonymous):

normal to one is normal to the other...

OpenStudy (anonymous):

okay,I got it :D Thank ya a lot.

OpenStudy (amistre64):

2 , 6 2k-2, 5k-4 ----------- 4k-4 + 30k-24 = 34k -28 ..... :)